# Towards Graph Foundation Models: A Study on the Generalization of Positional and Structural Encodings

### v1

@@ F.1 Convergence to PSEs Performance
This experiment is modeled after the experiment in Table 3 of Cantürk et al. [21]. Different downstream models (GCN, GatedGCN, GIN, and GPS) are trained using varying PSEs (none, rand, LapPE, RWSE, AllPSE) as well as GPSE, where as per usual we use a pre-trained GPSE node encoder trained on MolPCBA, on ZINC-12k. The exact architectures, like amount of layers and embedding sizes of the downstream models, vary based on previous studies. We refer to the human-readable YAML files in our code and the code of Cantürk et al. [21], for the exact architectures. As in Cantürk et al. [21], GCN, GatedGCN, and GIN are trained using Adam with a learning rate of , a weight decay of , for 1000 epochs with early stopping based on a validation split. Additionally, they use PyTorchs ReduceLROnPlateau to multiply the learning rate with 0.5 if the loss does not reduce for 10 consecutive epochs. GPS on other hand uses AdamW with the same learning rate and weight decay, but trained for 2000 epochs with the learning rate modified by PyTorchs CosineAnnealingWarmRestarts with 50 warmup epochs a typical choice for transformer models. For Table 14 we collect for each backbone using each PSE (except GPSE) the number of epochs until the best validation performance is achieved as well as the test performance at this point. We then compute how many epochs GPSE takes to reach this same test performance . We report that is the speedup of GPSE over .
F.2 Influence of Sample Size
@@
This experiment is modeled after the experiment in Table 3 and Table 4 of Cantürk et al. [21] considering only the GPS backbone. The GPS backbone is trained with varying PSEs (none, rand, LapPE, RWSE, AllPSE) as well as GPSE, GPSE-, and GPSE+on ZINC-12k. Also, a GINE backbone is trained with varying PSEs (LapPE, RWSE, AllPSE) as well as GPSE, GPSE-, and GPSE+on MolNet. All of these models are trained for a varying fraction of the dataset. The base case for this is a split of 80% for the training set, 10% for the validation set, and 10% for the test set. This 80% for the training set is then progressively reduced by factors of 2 without changing the size of the validation and test sets. The smallest training set fraction is for a total of 7 different training set fractions. We average each of these runs over 10 random seeds. Finally we subtract from each method except GPSE the performance of GPSE for the respective training set fraction and report the results in Fig. 2 and Fig. 3 including their standard deviation.
F.3 Generalization of PSEs
@@
This experiment is almost the same as Table 2. The only difference being that we do not report the speedup, but instead report the MAE on the test set in Table 16 as well as the train set in Table 5. In addition we also report the difference of the two in Table 3. Following the experiment from Fig. 2 and Fig. 3 we also conduct this experiment for the MolNet datasets training as before a GINE backbone instead of a GPS backbone for a slightly different set of PSEs. We present the AUROC on the test set in Table 17 and the difference between the AUROC on the test and train set in Table 4.
Appendix G Datasets
@@
MolPCBA [61] (MIT License) contains 400K small molecules derived from the MoleculeNet benchmark [68]. There are 323,555 unique molecular graphs in this dataset.
@@
ZINC-12k [69] (Custom license, free to use) is a 12K subset of the ZINC250K dataset [70]. Each graph is a molecule whose nodes are atoms (28 possible types) and whose edges are chemical bonds (3 possible types). The goal is to regress the constrained solubility [69] (logP) of the molecules. This dataset comes with a pre-defined split with 10K training, 1K validation, and 1K testing samples.
MoleculeNet small datasets [61] (MIT License) We follow Sun et al. [71] and use the selection of five small molecular property prediction datasets from the MoleculeNet benchmarks, including BBBP, BACE, Tox21, ToxCast, and SIDER. Each graph is a molecule, and it is processed the same way as for MolHIV and MolPCBA. All these datasets adopt the scaffold splitting strategy that is similarly used on MolHIV and MolPCBA.
CSL [69] (MIT License) contains 150 graphs that are known as circular skip-link graphs [49]. The goal is to classify each graph into one of ten isomorphism classes. The dataset is class-balanced, where each isomorphism class contains 15 graph instances. Splitting is done by stratified five-fold cross-validation.
EXP and CEXP [11] (unknown license) contain 600 pairs of graphs (1,200 graphs in total) that cannot be distinguished by 1&2-WL tests. The goal is to classify each graph into one of two isomorphism classes. Splitting is done by stratified five-fold cross-validation. CEXP is a modified version of EXP, where 50% of pairs are slightly modified to be distinguishable by 1-WL.
TRI and TRIX [12] (unknown license) contain 1000 3-regular 20-order graphs. The goal is to classify for each node whether it is contained in a triangle. Since the node colors of regular graphs are stable under 1-WL, MPNNs cannot tell nodes in these graphs apart and thus this task cannot be solved by vanilla MPNNs. TRIX is an extrapolation dataset of TRI, that shares the exact same training (and validation) dataset, however, the testset contains 1000 3-regular 100-order graphs. Splitting is done by five-fold cross-validation.
Appendix H GPSE Convergence to PSEs Performance
@@
Appendix I Influence of Sample Size on MolNet
@@
Appendix J Generalization of PSEs
Here we show the test performance in relation to Table 3 and Table 4. As already shown in [21], Table 16 and Table 17 show the test performance of various PSEs against various downstream backbones (on ZINC-12k showing MAE) and various MolNet datasets (showing AUROC). Notably, GPSE and its variants perform very well and usually either have the best performance or are only insignificantly worse than the best performance. Notably, for only 3 datasets (BACE, HIV, and MUV), some GPSE variants don’t belong to the best performers.

### v2

@@ E.1 Convergence to PSEs Performance
This experiment is modeled after the experiment in Table 3 of Cantürk et al. [2024]. Different downstream models (GCN, GatedGCN, GIN, and GPS) are trained using varying PSEs (none, rand, LapPE, RWSE, AllPSE) as well as GPSE, where as per usual we use a pre-trained GPSE node encoder trained on MolPCBA, on ZINC-12k. The exact architectures, like the number of layers and embedding sizes of the downstream models, vary based on previous studies. We refer to the human-readable YAML files in our code and the code of Cantürk et al. [2024], for the exact architectures. As in Cantürk et al. [2024], GCN, GatedGCN, and GIN are trained using Adam with a learning rate of , a weight decay of , for 1000 epochs with early stopping based on a validation split. Additionally, they use PyTorchs ReduceLROnPlateau to multiply the learning rate with 0.5 if the loss does not reduce for 10 consecutive epochs. GPS on other hand uses AdamW with the same learning rate and weight decay, but trained for 2000 epochs with the learning rate modified by PyTorchs CosineAnnealingWarmRestarts with 50 warmup epochs a typical choice for transformer models. For Table 14 we collect for each backbone using each PSE (except GPSE) the number of epochs until the best validation performance is achieved as well as the test performance at this point. We then compute how many epochs GPSE takes to reach this same test performance . We report that is the speedup of GPSE over .
E.2 Influence of Sample Size
@@
This experiment is modeled after the experiment in Table 3 and Table 4 of Cantürk et al. [2024] considering only the GPS backbone. The GPS backbone is trained with varying PSEs (none, rand, LapPE, RWSE, AllPSE) as well as GPSE, GPSE-, and GPSE+on ZINC-12k. Also, a GINE backbone is trained with varying PSEs (LapPE, RWSE, AllPSE) as well as GPSE, GPSE-, and GPSE+on MolNet. All of these models are trained for a varying fraction of the dataset. The base case for this is a split of 80% for the training set, 10% for the validation set, and 10% for the test set. This 80% for the training set is then progressively reduced by factors of 2 without changing the size of the validation and test sets. The smallest training set fraction is for a total of 7 different training set fractions. We average each of these runs over 10 random seeds. Finally, we subtract from each method except GPSE the performance of GPSE for the respective training set fraction and report the results in Fig. 2 and Fig. 3 including their standard deviation.
E.3 Generalization of PSEs
@@
This experiment is almost the same as Table 2. The only difference is that we do not report the speedup but instead report the MAE on the test set in Table 16 as well as the train set in Table 5. In addition, we also report the difference of the two in Table 3. Following the experiment from Fig. 2 and Fig. 3 we also conduct this experiment for the MolNet datasets training as before a GINE backbone instead of a GPS backbone for a slightly different set of PSEs. We present the AUROC on the test set in Table 17 and the difference between the AUROC on the test and train set in Table 4.
Appendix F Datasets
@@
MolPCBA [Hu et al., 2020] (MIT License) contains 400K small molecules derived from the MoleculeNet benchmark [Wu et al., 2017]. There are 323,555 unique molecular graphs in this dataset.
@@
ZINC-12k [Dwivedi et al., 2022c] (Custom license, free to use) is a 12K subset of the ZINC250K dataset [Gómez-Bombarelli et al., 2018]. Each graph is a molecule whose nodes are atoms (28 possible types) and whose edges are chemical bonds (3 possible types). The goal is to regress the constrained solubility [Dwivedi et al., 2022c] (logP) of the molecules. This dataset comes with a pre-defined split with 10K training, 1K validation, and 1K testing samples.
MoleculeNet small datasets [Hu et al., 2020] (MIT License) We follow Sun et al. [2022] and use the selection of five small molecular property prediction datasets from the MoleculeNet benchmarks, including BBBP, BACE, Tox21, ToxCast, and SIDER. Each graph is a molecule, and it is processed the same way as for MolHIV and MolPCBA. All these datasets adopt the scaffold splitting strategy that is similarly used on MolHIV and MolPCBA.
CSL [Dwivedi et al., 2022c] (MIT License) contains 150 graphs that are known as circular skip-link graphs [Murphy et al., 2019]. The goal is to classify each graph into one of ten isomorphism classes. The dataset is class-balanced, where each isomorphism class contains 15 graph instances. Splitting is done by stratified five-fold cross-validation.
EXP and CEXP [Abboud et al., 2021] (unknown license) contain 600 pairs of graphs (1,200 graphs in total) that cannot be distinguished by 1&2-WL tests. The goal is to classify each graph into one of two isomorphism classes. Splitting is done by stratified five-fold cross-validation. CEXP is a modified version of EXP, where 50% of pairs are slightly modified to be distinguishable by 1-WL.
TRI and TRIX [Sato et al., 2021] (unknown license) contain 1000 3-regular 20-order graphs. The goal is to classify for each node whether it is contained in a triangle. Since the node colors of regular graphs are stable under 1-WL, MPNNs cannot tell nodes in these graphs apart, and thus, this task cannot be solved by vanilla MPNNs. TRIX is an extrapolation dataset of TRI, that shares the exact same training (and validation) dataset, however, the testset contains 1000 3-regular 100-order graphs. Splitting is done by five-fold cross-validation.
Appendix G GPSE Convergence to PSEs Performance
@@
Appendix H Influence of Sample Size on MolNet
@@
Appendix I Generalization of PSEs
Here we show the test performance in relation to Table 3 and Table 4. As already shown in Cantürk et al. [2024], Table 16 and Table 17 show the test performance of various PSEs against various downstream backbones (on ZINC-12k showing MAE) and various MolNet datasets (showing AUROC). Notably, GPSE and its variants perform very well and usually either have the best performance or are only insignificantly worse than the best performance. Notably, for only 3 datasets (BACE, HIV, and MUV), some GPSE variants don’t belong to the best performers.